# OmniPred: Language Models as Universal Regressors

### v3

@@ 5 Experiments
    Is it possible to simultaneously regress on multiple tasks of different input spaces and objective scales?
@@
    Can finetuning improve accuracy over unseen studies outside of the pretraining set?
Appendix A contains additional ablations on OmniPred’s capabilities. Appendix B and C respectively contain details on language model and baseline implementations.
@@ 5.1 Simultaneous Regression
In Figure 4 for a model trained over real world data, we present an analogous visualization over hand-selected studies with drastically different input spaces, representative of objectives tuned internally at Google. These include standard machine learning (e.g. image classification and language modeling), production systems (e.g. Google bid simulation, LLM inference latency), and scientific research (e.g. protein and hardware design).
@@
Figure 4: Left: Diagonal fit (/) is better. Model’s prediction vs. ground truth objective over varying studies. Corporate-specific objective names are redacted. Right: Corresponding input spaces. “#-H, \$-T” is shorthand for a conditional hybrid input space with # root parameters and \$ total possible parameters.
@@ 5.2 Multi-task Transferrability
In this subsection, we demonstrate the model’s ability to transfer learn, i.e. improve accuracy over a specific task using knowledge gained from other similar but non-equivalent tasks, in contrast to “single-task” regressors (described in Appendix C) which only observe training data from the task being evaluated.
In Figure 5, we clearly see that the model’s accuracy improves with more tasks seen in training and eventually outperforms all traditional baselines. For AutoML studies encountered in Google, the error is averaged from a fixed subset of encountered studies. For BBOB, we can further demonstrate the model’s inter-study generalization capabilities over metadata $m$ (as opposed to $x$) by evaluating on unseen tasks with new shifts not encountered during training.
@@
In Figure 6, we further see that for the model, multi-task training consistently improves over single-task training, and in regimes with relatively lower input space saturation from training data, multi-task models outperform traditional baselines over several different domains. Interestingly, a single-task model trained from scratch remains a competitive choice and for certain domains such as AutoML, can even outperform all other single-task baselines. We hypothesize this is due to language-based representations being more appropriate for the conditional structures of these domains (e.g. AutoML).
@@
Vertex AI AutoML (Tabular) 1.4M 88 (4.6, 42.4)
Vertex AI AutoML (Text) 544K 118 56.0
Figure 6: Left: Lower ($\downarrow$) is better. Aggregate error across different domains. Right: Statistics on domains. Shorthand notation: “TpS” = Trials per Study, “SS” = Space Size, with brackets (#, \$) denoting conditional space with # root parameters and \$ total possible parameters.
5.3 Finetuning Analysis
@@
We first examine the conditions in which finetuning may be beneficial. In Table 5, we finetune various pretrained models over AutoML studies. While there is negligible benefit in finetuning the AutoML model on its data again, we see that a model pretrained over the entire Vizier dataset is able to finetune to the same level of accuracy as a pretrained AutoML model, while a BBOB-pretrained model leads to significantly worse results than even a single-task model. This suggests that knowledge obtained from pretraining can have a large (positive or negative) influence on transferrability over specific domains such as AutoML.
We further examine this effect by evaluating over unseen tasks, i.e. those which were newly created after the original training set was scraped, and can contain studies from new users and objectives. In Figure 7, we compare initialization from scratch (leading to single-task training) against a pretrained model on older Vizier data. We see that knowledge obtained from pretraining can significantly transfer over and help predictions over new tasks, although as seen on the left with three studies, there are few cases of negative transfer.
Method Mean Study Error ($\downarrow$)
@@
MLP Baseline 0.25
Tree Baseline 0.32
Random Forest 0.32
Gaussian Process 0.42
@@ 6 Conclusion
7 Limitations and Possible Extensions
Prompt-Side Numeric Tokenization: In this work, we directly represented numeric parameter values from $x$ into the default human readable format (e.g. $1234.5$ is serialized simply to ’1234.5’) to be consistent with LLM literature. This may be suboptimal, as the corresponding tokens may not exactly be digit-by-digit (e.g. SentencePiece tokenization leads to tokens {’12’, ’3’, ’4.5’}). One may instead potentially reuse the custom tokenization for $y$-values (e.g. <+><1><2><3><4><E0>) or in text-space, represent using other serializations which emphasize digits atomically, e.g. ’[1 10e2 2 10e1 3 10e0 4 10e-1 ]’) as in (Nogueira et al., 2021).
Pretrained English Encoder: Since $x$ includes parameter names and metadata which contain English words, warm-starting from a model pretrained on English text may improve accuracy. However, most checkpoints comparable to our model’s size (<1B params) such as T5-small and T5-medium are not pretrained over experimental data and are unlikely to understand the numerical meaning of e.g. ’learning_rate’. Furthermore, when using a pretrained English model, there are numerous confounding technical choices to consider (e.g. whether to freeze the encoder, tune the learning rate, embed additional custom float tokens, and use more English-based representations of $x$ and $m$), but this topic is worth pursuing in the future. In this work, we have already found that training a relatively small model from scratch can still achieve regression, thus suggesting our technique’s broad applicability even without English understanding.
Computational Costs: Compared to traditional baselines, a language model requires accelerator usage and has a relatively higher computational cost for both training and finetuning, in addition to higher inference times. In this work, we purposely designed the model to minimize costs by using $\approx$ 220M params which only requires at most 8 GPUs for training and 1 GPU for inference (see Appendix B).
Other Input Spaces: The Vizier API primarily focuses on hyperparameter tuning spaces. Traditionally, more complex spaces such as combinatorics and graphs require sophisticated modeling techniques to form regressors, largely in part to difficulties in representing the $x$’s as tensors. In addition, many applications with non-expressible spaces such program synthesis are impossible to traditionally regress over. We believe that text and token-based representations are highly promising and are widely applicable to domains previously unexplored in the field of experimental design.
Acknowledgements
We would like to thank Olivier Bachem, Hado van Hasselt, John Jumper, Aviral Kumar, Yingjie Miao, Sebastian Nowozin, Mangpo Phothilimthana, Zi Wang, Scott Yak, and Amir Yazdanbakhsh for useful discussions and Daniel Golovin for continuing support.
@@ Acknowledgements
Appendix A Model Ablations
We ablate below certain settings and scenarios which affect the model’s prediction accuracy.
@@ A.1 $y$-Tokenization
A.2 Effect of Sampling
The LM can output extreme outliers in its $y$-prediction, usually due to an inaccurate prediction on the exponent token or significant digits. While such issues do not occur once the model has nearly perfectly regressed on a task (e.g. BBOB), they occur frequently on realistic tasks with nontrivial error (e.g. AutoML) and thus require techniques for correction.
One obvious method to reduce error is to increase sample count:
We further compare aggregation methods and see that using the median considerably outperforms both max-likelihood and mean. We hypothesize this is due to the median’s robustness to hallucinated outlier samples which can occur with relatively high probability and can also skew the mean.
Mean Study Error ($\downarrow$)
Empirical Aggregation Method AutoML (Full 540K) BBOB (Full 1M)
Median (default) 0.15 0.01
Max-likelihood 0.22 0.01
Mean 0.23 0.01
Table 7: Lower ($\downarrow$) is better. Comparisons between different sample aggregation methods when using 64 samples.
A.3 Uncertainty
We begin by demonstrating the LM’s ability to nonparametrically express multimodal distributions in Figure 9 when we train the model against randomly sign-flipped BBOB objectives. In contrast, traditional methods such as ensembled MLPs (Havasi et al., 2021) and Gaussian Process mixtures (Bonilla et al., 2007) must specify the mixture count a priori.
Figure 9: Setting similar to Figure 3 in the main section of the paper, but with bimodality.
Pearson, Kendall-Tau, Spearman Correlation ($\uparrow$)
Regressor Uncertainty Metric AutoML BBOB
A.4 Ranking and Correlation Metrics
@@
Regressor BBOB Bid Simulation Google-AML Init2Winit Protein Design V-AML (Tab) V-AML (Text)
@@ Appendix C Baseline Details
C.1 Vizier Input Space
@@ C.4 Regressor Baselines
Below, we list out the specific implementation details of our regressor baselines. One nuanced issue is of hyperparameter tuning the regressors themselves, which could affect results. In order to be reasonably fair to all regressors (including our own OmniPred which has its own hyperparameters), for each regressor, we used a reasonable fixed set of hyperparameters for consistent throughout all experiments.
@@
Gaussian Process: The GP regressor model is from the GP-Bandit implementation found Open Source Vizier at https://github.com/google/vizier and consists of the following:
  • $\alpha\sim\mathrm{TruncatedLogNormal}$ controls the amplitude of Matern5/2 kernel.
@@
Tree and Random Forest: We use the standard API (XGBRegressor, XGBRFRegressor) found in XGBoost (https://github.com/dmlc/xgboost) with default hyperparameters. Although tree-based methods do not generally require rescaling, we still applied consistent $x$-preprocessing (in particular to deal with optional log or reverse-log scaling).
@@ D.1 Study Preprocessing
Filtering specific users: There are specific human and automated “power users” which produce orders of magnitude more studies than the average user. Some automated users in particular are simply automatic unit tests involving the use of Vizier. We disregard studies from these users to prevent them from dominating the data distribution.
@@ D.2 Real World Data Descriptions
Bid Simulation: Contains hyperparameters for one of Google’s bid simulators. The simulator estimates how advertisements might have performed in terms of key metrics such as cost, impressions, clicks, and conversion volume.
Google AutoML (Internal): A Google internal version of Vertex AI AutoML. Uses a different input space and a different ML pipeline implementation than Vertex AI AutoML.
Init2Winit: A Google research project for running deterministic, scalable, and well-documented deep learning experiments, with a particular emphasis on optimization and tuning experiments (e.g. ResNets on CIFAR10, Transformers on LM1B). Public codebase can be found in https://github.com/google/init2winit.
@@
Vertex AI AutoML (Tabular and Text): A Vertex AI platform for automated ML model selection and training for tabular or text data. For tabular data, AutoML searches over a tree of model and optimizer types, their hyperparameters, data transformation, and other components in the ML pipeline. For text, AutoML trains an ML model to classify text data, extract information, or understand the sentiment of the authors. See https://cloud.google.com/vertex-ai?#train-models-with-minimal-ml-expertise.
@@ D.3 Serialization Examples
For transparency, we provide examples of text representations seen by the model. Disclaimer: Due to corporate privacy policies, we redacted (in red) some parameter names and values.
@@
Google AutoML batch_size: 128 model_type: REDACTEDactivation_fn: "tanh"batch_norm: "True"dropout: 0.143embedding_combiner: "mean"gradient_clip_norm: 1.63e+03num_hidden_layers: 1hidden_units[0]: 359optimizer_type: "AdamOptimizer"beta1: 0.9beta2: 0.999learning_rate: 0.926nlp_vocabulary_strategy:     "adjusted_mutual_info"vocabulary_strategy:     "adjusted_mutual_info" title: "n-w597ng99i7lj0-q40zcboi1ea7l" user: REDACTEDdescription: ""objective: "val_categorical_cross_entropy"amc_model_version: REDACTEDtask_type: "multi_class_classification"

### v4

@@ 5 Experiments
    Can we simultaneously regress on multiple tasks of different input spaces and objective scales?
@@
    Does online finetuning improve accuracy over unseen studies outside of the pretraining set?
@@ 5.1 Simultaneous Regression
Figure 4: Left: Diagonal fit (/) is better. Model’s $y$-prediction vs. ground truth over varying studies. Corporate-specific objective names are redacted. Right: Corresponding input spaces. “#-H, \$-T” is shorthand for a conditional hybrid input space with # root parameters and \$ total possible parameters.
In Figure 4 for a model trained over real world data, we present an analogous visualization over hand-selected studies with drastically different input spaces, representative of objectives tuned commonly in real world settings. These include standard machine learning (e.g. image classification and language modeling), production systems (e.g. bid simulation, LLM inference latency), and scientific research (e.g. protein and hardware design).
@@ 5.2 Multi-task Transferrability
In this subsection, we demonstrate the model’s ability to transfer learn, i.e. improve accuracy over a specific task using knowledge gained from other similar but non-equivalent tasks, in contrast to “single-task” regressors (described in Appendix C) which only observe training data from the task being evaluated. Note that single-task baselines such as MLPs are incapable of multi-task training when tasks have different dimensionalities.
In Figure 5, we clearly see that the model’s accuracy improves with more tasks seen in training and eventually outperforms all traditional baselines. For AutoML studies, the error is averaged from a fixed subset of encountered studies. For BBOB, we can further demonstrate the model’s inter-study generalization capabilities over metadata $m$ (as opposed to $x$) by evaluating on unseen tasks with new shifts not encountered during training.
@@
In Figure 6, we further see that for the model, multi-task training consistently improves over single-task training, and in regimes with relatively lower input space saturation (i.e. low trial to parameter count ratio) from training data, multi-task models outperform traditional baselines over several different domains. Interestingly, a single-task model trained from scratch remains a competitive choice and for certain domains such as AutoML, can even outperform all other single-task baselines. We hypothesize this is due to language-based representations being more appropriate for the conditional structures of these domains (e.g. AutoML).
@@
Vertex AI (Tabular) 1.4M 88 (4.6, 42.4)
Vertex AI (Text) 544K 118 56.0
Figure 6: Left: Lower ($\downarrow$) is better. Aggregate error across different domains. Right: Statistics on domains. Shorthand notation: “TpS” = Trials per Study, “SS” = Space Size, with brackets (#, \$) denoting conditional space with # root parameters and \$ total possible parameters. Note that all baselines are single-task regressors.
5.3 Online Finetuning Analysis
@@
We first examine the conditions in which finetuning may be beneficial. In Table 5, we finetune various pretrained models over AutoML studies. While there is negligible benefit in finetuning the AutoML model on its data again, we see that a model pretrained over the entire real world dataset is able to finetune to the same level of accuracy as a pretrained AutoML model, while a BBOB-pretrained model leads to significantly worse results than even a single-task model. This suggests that knowledge obtained from pretraining can have a large (positive or negative) influence on transferrability over specific domains such as AutoML.
Method Mean Study Err. ($\downarrow$)
@@
MLP (Baseline) 0.25
Tree (Baseline) 0.32
Random Forest (Baseline) 0.32
Gaussian Process (Baseline) 0.42
@@
We further examine this effect by evaluating over unseen tasks, i.e. those which were newly created after the original training set was scraped, and can contain studies from new users and objectives. In Figure 7, we compare initialization from scratch (leading to single-task training) against a pretrained model on older real world data. We see that knowledge obtained from pretraining can significantly transfer over and help predictions over new tasks, although as seen on the left with three studies, there are few cases of negative transfer.
6 Ablations
We further ablate certain important settings and scenarios which affect the model’s prediction accuracy below. Appendix A contains additional ablations, specifically (1) around our choice of $y$-tokenization and (2) experimental outcomes when using ranking-based regression metrics.
6.1 Effect of Sampling
The LM can output extreme outliers in its $y$-prediction, usually due to an inaccurate prediction on the exponent token or significant digits. While such issues do not occur once the model has nearly perfectly regressed on a task (e.g. BBOB), they occur frequently on realistic tasks with nontrivial error (e.g. AutoML) and thus require techniques for correction. One obvious method to reduce error is to increase sample count, as demonstrated in Figure 8.
An additional method is to change the method of aggregation across samples. In Table 6, we see that using the median considerably outperforms both max-likelihood and mean. We hypothesize this is due to the median’s robustness to hallucinated outlier samples which can occur with relatively high probability and can also skew the mean.
Mean Study Error ($\downarrow$)
Empirical Aggregation Method AutoML (Full 540K) BBOB (Full 1M)
Median (default) 0.15 0.01
Max-likelihood 0.22 0.01
Mean 0.23 0.01
Table 6: Lower ($\downarrow$) is better. Comparisons between different aggregation methods when using 64 samples when using full datasets for pretraining.
6.2 Uncertainty Calibration
Figure 9: Setting similar to Figure 3, but with bimodality.
We begin by demonstrating the LM’s ability to nonparametrically express distributions with multiple modes in Figure 9 when we train the model against randomly sign-flipped BBOB objectives. In contrast, traditional methods such as ensembled MLPs (Havasi et al., 2021) and Gaussian Process mixtures (Bonilla et al., 2007) must specify the mixture count a priori.
Pearson, Kendall-Tau, Spearman
Regressor Uncertainty Metric AutoML BBOB
6.3 Study Size vs. Multi-task Gain
7 Discussion: Limitations and Extensions
In-Context Learning (ICL): By design, we maximized the allowed prompt length to allow flexibility in representing $(x,m)$ and thus did not use in-context regression, since a single $(x,m)$ prompt could be 10K+ tokens long for some applications. While online finetuning in principle allows infinite amounts of data to be absorbed at inference time, it is worth investigating ICL methods which allow arbitrarily long prompts as well. Current methods (Chen et al., 2022) require significant input compression are which only applicable to tabular-like $x$’s. Additional use of ChatGPT and other service-based chat APIs (Vacareanu et al., 2024) rely on regression as an emergent property of expensive LLM-based training, making their serious use difficult, especially as such services cannot absorb large amount of user-defined offline $(x,y)$ data which often contain the most relevant information for finetuning.
Prompt-Side Numeric Tokenization: In this work, we directly represented numeric parameter values from $x$ into the default human readable format (e.g. $1234.5$ is serialized simply to ‘1234.5’) to be consistent with LLM literature. This may be suboptimal, as the corresponding tokens may not exactly be digit-by-digit (e.g. T5’s tokenization leads to tokens {‘12’, ‘3’, ‘4.5’}). One may instead potentially reuse the custom tokenization for $y$-values (e.g. <+><1><2><3><4><E0>) or in text-space, represent using other serializations which emphasize digits atomically, e.g. ‘[1 10e2 2 10e1 3 10e0 4 10e-1 ]’) as in (Nogueira et al., 2021).
Pretrained Language Encoder: Since $x$ includes parameter names and metadata which contain English words, warm-starting from a model pretrained on English text may improve accuracy. However, most checkpoints comparable to our model’s size (<1B params) are not pretrained over experimental data and are unlikely to understand the numerical meaning of e.g. ‘learning_rate’. Furthermore, using a pretrained English model introduces numerous confounding technical choices to consider (e.g. whether to freeze the encoder, tune the learning rate, embed additional custom float tokens, and use more English-based representations of $x$ and $m$), but this topic is worth pursuing in the future. In this work, it is already surprising that a relatively small model trained from scratch can still achieve regression, thus suggesting our technique’s broad applicability even without English understanding.
Computational Costs: Compared to traditional baselines, a language model requires accelerator usage and has a relatively higher computational cost for both training and finetuning, in addition to higher inference times. In this work, we purposely designed the model to minimize costs by using $\approx$200M params which only requires at most 8 GPUs for training and 1 GPU for inference (see Appendix B).
Other Input Spaces: The OSS Vizier API primarily focuses on hyperparameter tuning spaces. Traditionally, more complex spaces such as combinatorics and graphs require sophisticated modeling techniques to form regressors, largely in part to difficulties in representing the $x$’s as tensors. In addition, many applications with non-expressible spaces such program synthesis are impossible to traditionally regress over. We believe that text and token-based representations are highly promising and are widely applicable to domains previously unexplored in the field of experimental design.
8 Impact Statement
This research addresses the ability to regress metrics against textual data. Since any textual metadata may be collected, user-specific information may be used, which raises privacy concerns. This would be particularly true on sensitive topics (e.g. predicting metrics related to personal protected characteristics).
Our research does not involve such sensitive topics for regression, as it is performed over blackbox optimization data. We followed ethical guidelines as all users in our proprietary dataset have consented to have their tuning data saved and used for offline analysis. The proprietary real world dataset does not contain any sensitive personal information other than usernames. Furthermore, we do not released any of the trained checkpoints, as it may be possible to reverse-engineer parts of the training data, which can lead privacy violations and data leakage.
@@ 9 Conclusion
Appendix A Additional Ablations
We further ablate certain settings and scenarios which affect the model’s prediction accuracy below.
@@ A.1 $y$-Tokenization
A.2 Ranking and Correlation Metrics
@@
Regressor BBOB Bid Simulation AutoML Init2Winit Protein Design V-AI (Tab) V-AI (Text)
@@ Appendix C Baseline Details
C.1 OSS Vizier Input Space
@@ C.4 Regressor Baselines
Below, we list out the specific implementation details of our regressor baselines. One nuanced issue is of hyperparameter tuning the regressors themselves, which could affect results. In order to be reasonably fair to all regressors (including our own OmniPred which has its own hyperparameters), for each regressor, we used a reasonable fixed set of hyperparameters for consistency throughout all experiments.
@@
Gaussian Process: The GP regressor model is from the GP-Bandit implementation (Song et al., 2024) found in Open Source Vizier at https://github.com/google/vizier and consists of the following:
  • $\alpha\sim\mathrm{TruncatedLogNormal}$ controls the amplitude of Matern-5/2 kernel.
@@
Tree and Random Forest: We use the standard API (XGBRegressor, XGBRFRegressor in https://github.com/dmlc/xgboost) found in XGBoost (Chen & Guestrin, 2016), with the following grid-search hyperparameter sweeps for each study:
  • ‘‘min_child_weight": [1, 5, 10]
  • ‘‘learning_rate": [0.001, 0.01, 0.1]
  • ‘‘gamma": [0.0, 0.3, 0.5]
  • ‘‘subsample": [0.6, 0.8, 1.0]
  • ‘‘colsample_bytree": [0.6, 0.8, 1.0]
  • ‘‘max_depth": [3, 5, 7]
Although tree-based methods do not generally require rescaling, we still applied consistent $x$-preprocessing (in particular to deal with optional log or reverse-log scaling).
@@ D.1 Study Preprocessing
Filtering specific users: There are specific human and automated “power users” which produce orders of magnitude more studies than the average user. Some automated users in particular are simply automatic unit tests involving the use of OSS Vizier. We disregard studies from these users to prevent them from dominating the data distribution.
@@ D.2 Real World Data Descriptions
Bid Simulation: Contains hyperparameters for proprietary bid simulation. The simulator estimates how advertisements might have performed in terms of key metrics such as cost, impressions, clicks, and conversion volume.
AutoML: Collection of proprietary AutoML data for tuning user objectives.
Init2Winit: Data from running deterministic, scalable, and well-documented deep learning experiments, with a particular emphasis on optimization and tuning experiments (e.g. ResNets on CIFAR10, Transformers on LM1B). Public codebase can be found in https://github.com/google/init2winit.
@@
Vertex AI (Tabular and Text): A Vertex AI platform for automated ML model selection and training for tabular or text data. For tabular data, Vertex AI searches over a tree of model and optimizer types, their hyperparameters, data transformation, and other components in the ML pipeline. For text, Vertex AI trains an ML model to classify text data, extract information, or understand the sentiment of the authors. For more information, see:
https://cloud.google.com/vertex-ai?#train-models-with-minimal-ml-expertise.
@@ D.3 Serialization Examples
For transparency, we provide examples of text representations seen by the model. Disclaimer: Due to privacy policies, we redacted (in red) some parameter names and values.
@@
AutoML batch_size: 128 model_type: REDACTED activation_fn: "tanh" batch_norm: "True" dropout: 0.143 embedding_combiner: "mean" gradient_clip_norm: 1.63e+03 num_hidden_layers: 1 hidden_units[0]: 359 optimizer_type: "AdamOptimizer" beta1: 0.9 beta2: 0.999 learning_rate: 0.926 nlp_vocabulary_strategy:     "adjusted_mutual_info" vocabulary_strategy:     "adjusted_mutual_info" title: "n-w597ng99i7lj0-q40zcboi1ea7l" user: REDACTED description: "" objective: "val_categorical_cross_entropy" amc_model_version: REDACTED task_type: "multi_class_classification"